OpenStudy (anonymous):

Use Newton's method with the specified initial approximation x 1 to find x 3, the third approximation to the root of the given equation. (Round your answer to four decimal places). x^3 + 4x − 3 = 0, x 1 = 1

OpenStudy (zale101):

Newton's method: \(x_2=x_1-\Large\frac{f(x_1)}{f'(x_1)}\) \(x_3=x_2-\Large\frac{f(x_2)}{f'(x_2)}\) and so on. Makes sense? So you have x^3 + 4x − 3 = 0, x 1 = 1

OpenStudy (anonymous):

okay I understand that part

OpenStudy (zale101):

to find \(x_2\), you must take whatever the number of \(x_1\), which is 1, subtract it by the f(1), divided by the derivative f'(1).

OpenStudy (zale101):

So \(f(x)=x^3 + 4x − 3 \) and \(x_1=1\) \(x_2=1-\Large\frac{f(1)}{f'(1)}\)

OpenStudy (dan815):

Hi would you like a more detailed explanation of newtons method